You have plenty of cash to invest. You are considering an investment of $125,000 in a project which is expected to earn $14425 a year for ten years. Is it an attractive investment if your minimum expected annual rate of return is 5% (compound interest)? Calculate the expected annual rate of return (R) of the project. (6 points) NO TABLE and use a simple equation. No hit or miss trial
Answer #1

ANSWER:

PV = Cash flow per period * ( ( 1 - (1 + i) ^ - n) / i )

cash flow per period = 14,425

pv = 125,000

n = 10

i = ?

125,000 = 14,425 * ( ( 1 - (1 + i) ^ - 10) / i )

125,000 / 14,425 = (1 - (1 + i) ^ - 10 ) / i )

8.6655 = (1 - (1 + i) ^ - 10 ) / i )

8.6655 * i = 1 - (1 + i) ^ - 10

8.6655 i = 1 / (1 + i ) ^ 10

8.6655 i * (1 + i) ^ 10 = 1

i * (1 + i) ^ 10 = 1 / 8.6655

i * (1 + i) ^ 10 = 0.1154

(1 + i) ^ 10 = 0.1154 / i

1 + i = (0.1154 / i) ^ 1 / 10

1 + i = 0.8057 / (i) ^ 1 / 10

i = 0.8057 / (i) ^ 1 / 10 - 1

i = 0.0269 or 2.69%

so the annual rate of return is 2.69%

since the annual rate of return is less then the marr , therefore the investment is not attractive. (2.69% < 5%)